# Methodological Insights into Structural Causal Modelling and Uncertainty-Aware Forecasting for Economic Indicators

### v1

@@ Abstract
This paper presents a methodological approach to financial time series analysis by combining causal discovery and uncertainty-aware forecasting. As a case study, we focus on four key U.S. macroeconomic indicators – GDP, economic growth, inflation, and unemployment – and we apply the LPCMCI framework with Gaussian Process Distance Correlation (GPDC) to uncover dynamic causal relationships in quarterly data from 1970 to 2021. Our results reveal a robust unidirectional causal link from economic growth to GDP and highlight the limited connectivity of inflation, suggesting the influence of latent factors. Unemployment exhibits strong autoregressive dependence, motivating its use as a case study for probabilistic forecasting. Leveraging the Chronos framework, a large language model trained for time series, we perform zero-shot predictions on unemployment. This approach delivers accurate forecasts one and two quarters ahead, without requiring task-specific training. Crucially, the model’s uncertainty-aware predictions yield 90% confidence intervals, enabling effective anomaly detection through statistically principled deviation analysis. This study demonstrates the value of combining causal structure learning with probabilistic language models to inform economic policy and enhance forecasting robustness.
@@ 1 Introduction
Time series analysis is foundational in financial modelling, supporting tasks ranging from risk management and portfolio optimisation to macroeconomic forecasting. In multivariate financial settings – such as analysing interest rates, exchange rates, asset prices, and credit indicators simultaneously – uncovering the causal relationships among variables is crucial for both interpretability and intervention. Structural causal models for time series, such as the LPCMCI framework (see Section˜2.1), enable the identification of dynamic and contemporaneous causal links through conditional independence testing, offering a principled approach to disentangle complex interdependencies in financial systems. On the other hand, forecasting a single financial variable – such as a stock price, inflation rate, or cryptocurrency valuation – under uncertainty is equally vital. Recent advances in large pre-trained models based on language model architectures, such as Chronos [1] (see Section˜2.2), [^1] https://github.com/amazon-science/chronos-forecasting (on 23 May 2025). have enabled the generation of probabilistic forecasts that capture the full distribution of possible outcomes. This is particularly valuable in finance, where understanding the range and likelihood of future scenarios directly informs trading strategies, risk exposure assessments, and regulatory compliance.
In this study, we analyse the causal interplay among four key macroeconomic indicators in the United States – Gross Domestic Product (GDP), economic growth, inflation, and unemployment – over the period 1970 to 2021. Consistent with the methodology proposed by [7], we employ quarterly aggregated data, normalised via standard scaling to ensure numerical stability and comparability across variables. The resulting causal graph (see Section˜4.1) highlights a robust unidirectional link from Economic Growth to GDP, with no evidence of the reverse, while Inflation remains weakly connected and potentially influenced by latent factors. In contrast, Unemployment exhibits strong autoregressive structure, justifying its selection for targeted probabilistic forecasting.
Given the pronounced temporal self-dependence observed in the Unemployment series – characterised by robust causal links to its own lagged values across consecutive quarters – we focus the subsequent analysis on this variable. Specifically, we investigate the predictive structure of Unemployment using a zero-shot probabilistic forecasting approach based on a large pretrained language model.
@@ 3 Data Selection and Exploration
Following [7], we conducted a structural causal analysis on the interrelationships among key macroeconomic indicators in the United States – namely GDP, inflation, economic growth, and unemployment – over the period from 1970 to 2021, using quarterly normalised data. To facilitate comparison and improve the numerical stability of the causal analysis, as recommended by [7], we normalised all variables using standard scaling. This transformation preserves the underlying structure and relationships among variables, making it especially suitable for causal discovery in multivariate time series.
@@ 4.1 Causal Analysis with GPDC
More unexpected, however, is the behaviour of inflation. In Figure˜4, inflation appears to be only loosely connected with the other macroeconomic indicators – and even lacks a strong autoregressive component. This weak integration may indicate that the temporal evolution of inflation is governed by latent or exogenous factors not captured in the present model, such as global commodity prices, monetary policy shocks, or anticipatory expectations not directly observable in standard macroeconomic aggregates. In contrast, the variable Unemployment displays a markedly self-referential pattern, with strong quarter-over-quarter dependencies suggesting high temporal inertia. This observation aligns with known frictions in labour markets, such as delayed hiring/firing responses, regulatory constraints, or mismatch between labour supply and demand.
Given the pronounced temporal self-dependence observed in the Unemployment series – characterised by robust causal links to its own lagged values across consecutive quarters – we investigate the predictive structure of Unemployment using a zero-shot probabilistic forecasting approach based on a large pretrained language model.
@@ 5 Zero-Shot Uncertainty-Aware Time Series Prediction
Crucially, this probabilistic treatment allows us to represent the total predictive uncertainty, which encompasses both aleatory and epistemic components. Aleatory uncertainty reflects inherent variability in the data – such as stochastic fluctuations in unemployment rates due to unforeseen macroeconomic shocks – and cannot be reduced by additional observations. Epistemic uncertainty, by contrast, arises from limited knowledge or model uncertainty, and is particularly relevant in zero-shot or distribution-shift settings where the model has not encountered similar patterns during pretraining. While Chronos does not explicitly disentangle these sources, its predictive intervals widen in regions of higher overall uncertainty, which may stem from either or both components. This aligns with recent efforts in developing uncertainty-aware deep learning models [5, 8, 12, 13, 2, 3], aimed at improving decision-making reliability in high-stakes domains.
@@
Regarding calibration, the 90% prediction intervals produced by Chronos cover approximately 81% of the 42 forecasted observations (about 8 violations). While this undercoverage suggests some degree of miscalibration, it is important to note that (i) the evaluation set is small, so sampling variability is high – under perfect calibration, the number of violations follows a $\text{Binomial}(42,0.1)$ distribution, where observing about 8 violations is within two standard deviations of the expected value; and (ii) the largest errors coincide with unprecedented shocks, where both aleatory and epistemic uncertainties are elevated. Chronos does not guarantee exact nominal coverage, but its intervals adaptively widen after shocks, reflecting increased total predictive uncertainty.
@@ 6 Conclusion and Future Work
This study demonstrates the complementary strengths of structural causal models and uncertainty-aware forecasting in analysing macroeconomic time series. By applying the LPCMCI framework combined with the GPDC conditional independence test, we uncover meaningful causal relations among GDP, economic growth, inflation, and unemployment – revealing, for instance, the autoregressive dynamics of unemployment. Building on this, we employ a zero-shot probabilistic forecasting approach using Chronos to model unemployment, showcasing the feasibility of high-accuracy, uncertainty-calibrated predictions without retraining. Our results underscore the potential of integrating causal discovery with uncertainty-aware forecasting to support robust data interpretation and policy design. The capacity to flag anomalies through distributional deviation further enhances the practical utility of such models.

### v2

@@ Abstract
This paper presents a methodological approach to financial time series analysis by combining causal discovery and uncertainty-aware forecasting. As a case study, we focus on four key U.S. macroeconomic indicators  — GDP, economic growth, inflation, and unemployment  — and we apply the LPCMCI framework with Gaussian Process Distance Correlation (GPDC) to uncover dynamic causal relationships in quarterly data from 1970 to 2021. Our results reveal a robust unidirectional causal link from economic growth to GDP and highlight the limited connectivity of inflation, suggesting the influence of latent factors. Unemployment exhibits strong autoregressive dependence, motivating its use as a case study for probabilistic forecasting. Leveraging the Chronos framework, a large language model trained for time series, we perform zero-shot predictions on unemployment. This approach delivers accurate forecasts one and two quarters ahead, without requiring task-specific training. Crucially, the model’s uncertainty-aware predictions yield 90% confidence intervals, enabling effective anomaly detection through statistically principled deviation analysis. This study demonstrates the value of combining causal structure learning with probabilistic language models to inform economic policy and enhance forecasting robustness.
@@ 1 Introduction
Time series analysis is foundational in financial modelling, supporting tasks ranging from risk management and portfolio optimisation to macroeconomic forecasting. In multivariate financial settings  — such as analysing interest rates, exchange rates, asset prices, and credit indicators simultaneously  — uncovering the causal relationships among variables is crucial for both interpretability and intervention. Structural causal models for time series, such as the LPCMCI framework (see Section˜2.1), enable the identification of dynamic and contemporaneous causal links through conditional independence testing, offering a principled approach to disentangle complex interdependencies in financial systems. On the other hand, forecasting a single financial variable  — such as a stock price, inflation rate, or cryptocurrency valuation  — under uncertainty is equally vital. Recent advances in large pre-trained models based on language model architectures, such as Chronos [1] (see Section˜2.2), [^1] https://github.com/amazon-science/chronos-forecasting (on 23 May 2025). have enabled the generation of probabilistic forecasts that capture the full distribution of possible outcomes. This is particularly valuable in finance, where understanding the range and likelihood of future scenarios directly informs trading strategies, risk exposure assessments, and regulatory compliance.
In this study, we analyse the causal interplay among four key macroeconomic indicators in the United States  — Gross Domestic Product (GDP), economic growth, inflation, and unemployment  — over the period 1970 to 2021. Consistent with the methodology proposed by [7], we employ quarterly aggregated data, normalised via standard scaling to ensure numerical stability and comparability across variables. The resulting causal graph (see Section˜4.1) highlights a robust unidirectional link from Economic Growth to GDP, with no evidence of the reverse, while Inflation remains weakly connected and potentially influenced by latent factors. In contrast, Unemployment exhibits strong autoregressive structure, justifying its selection for targeted probabilistic forecasting.
Given the pronounced temporal self-dependence observed in the Unemployment series  — characterised by robust causal links to its own lagged values across consecutive quarters  — we focus the subsequent analysis on this variable. Specifically, we investigate the predictive structure of Unemployment using a zero-shot probabilistic forecasting approach based on a large pretrained language model.
@@ 3 Data Selection and Exploration
Following [7], we conducted a structural causal analysis on the interrelationships among key macroeconomic indicators in the United States  — namely GDP, inflation, economic growth, and unemployment  — over the period from 1970 to 2021, using quarterly normalised data. To facilitate comparison and improve the numerical stability of the causal analysis, as recommended by [7], we normalised all variables using standard scaling. This transformation preserves the underlying structure and relationships among variables, making it especially suitable for causal discovery in multivariate time series.
@@ 4.1 Causal Analysis with GPDC
More unexpected, however, is the behaviour of inflation. In Figure˜4, inflation appears to be only loosely connected with the other macroeconomic indicators  — and even lacks a strong autoregressive component. This weak integration may indicate that the temporal evolution of inflation is governed by latent or exogenous factors not captured in the present model, such as global commodity prices, monetary policy shocks, or anticipatory expectations not directly observable in standard macroeconomic aggregates. In contrast, the variable Unemployment displays a markedly self-referential pattern, with strong quarter-over-quarter dependencies suggesting high temporal inertia. This observation aligns with known frictions in labour markets, such as delayed hiring/firing responses, regulatory constraints, or mismatch between labour supply and demand.
Given the pronounced temporal self-dependence observed in the Unemployment series  — characterised by robust causal links to its own lagged values across consecutive quarters  — we investigate the predictive structure of Unemployment using a zero-shot probabilistic forecasting approach based on a large pretrained language model.
@@ 5 Zero-Shot Uncertainty-Aware Time Series Prediction
Crucially, this probabilistic treatment allows us to represent the total predictive uncertainty, which encompasses both aleatory and epistemic components. Aleatory uncertainty reflects inherent variability in the data  — such as stochastic fluctuations in unemployment rates due to unforeseen macroeconomic shocks  — and cannot be reduced by additional observations. Epistemic uncertainty, by contrast, arises from limited knowledge or model uncertainty, and is particularly relevant in zero-shot or distribution-shift settings where the model has not encountered similar patterns during pretraining. While Chronos does not explicitly disentangle these sources, its predictive intervals widen in regions of higher overall uncertainty, which may stem from either or both components. This aligns with recent efforts in developing uncertainty-aware deep learning models [5, 8, 12, 13, 2, 3], aimed at improving decision-making reliability in high-stakes domains.
@@
Regarding calibration, the 90% prediction intervals produced by Chronos cover approximately 81% of the 42 forecasted observations (about 8 violations). While this undercoverage suggests some degree of miscalibration, it is important to note that (i) the evaluation set is small, so sampling variability is high  — under perfect calibration, the number of violations follows a $\text{Binomial}(42,0.1)$ distribution, where observing about 8 violations is within two standard deviations of the expected value; and (ii) the largest errors coincide with unprecedented shocks, where both aleatory and epistemic uncertainties are elevated. Chronos does not guarantee exact nominal coverage, but its intervals adaptively widen after shocks, reflecting increased total predictive uncertainty.
@@ 6 Conclusion and Future Work
This study demonstrates the complementary strengths of structural causal models and uncertainty-aware forecasting in analysing macroeconomic time series. By applying the LPCMCI framework combined with the GPDC conditional independence test, we uncover meaningful causal relations among GDP, economic growth, inflation, and unemployment  — revealing, for instance, the autoregressive dynamics of unemployment. Building on this, we employ a zero-shot probabilistic forecasting approach using Chronos to model unemployment, showcasing the feasibility of high-accuracy, uncertainty-calibrated predictions without retraining. Our results underscore the potential of integrating causal discovery with uncertainty-aware forecasting to support robust data interpretation and policy design. The capacity to flag anomalies through distributional deviation further enhances the practical utility of such models.